# SEA-SQL: Semantic-Enhanced Text-to-SQL with Adaptive Refinement

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly contributed to the progress of the Text-to-SQL task. A common requirement in many of these works is the post-correction of SQL queries. However, the majority of this process entails analyzing error cases to develop prompts with rules that eliminate model bias. And there is an absence of execution verification for SQL queries. In addition, the prevalent techniques primarily depend on GPT-4 and few-shot prompts, resulting in expensive costs. To investigate the effective methods for SQL refinement in a cost-efficient manner, we introduce Semantic-Enhanced Text-to-SQL with Adaptive Refinement (SEA-SQL), which includes Adaptive Bias Elimination and Dynamic Execution Adjustment, aims to improve performance while minimizing resource expenditure with zero-shot prompts. Specifically, SEA-SQL employs a semantic-enhanced schema to augment database information and optimize SQL queries. During the SQL query generation, a fine-tuned adaptive bias eliminator is applied to mitigate inherent biases caused by the LLM. The dynamic execution adjustment is utilized to guarantee the executability of the bias eliminated SQL query. We conduct experiments on the Spider and BIRD datasets to demonstrate the effectiveness of this framework. The results demonstrate that SEA-SQL achieves state-of-the-art performance in the GPT3.5 scenario with 9%-58% of the generation cost. Furthermore, SEA-SQL is comparable to GPT-4 with only 0.9%-5.3% of the generation cost.
@@ 1 Introduction
Relational databases play a pivotal role in information technology by offering a structured and organized framework for data storage. Through tables, rows, and columns, data is efficiently organized, ensuring clarity and ease of management. However, users typically interact with databases through SQL queries, which is a barrier for those unfamiliar with SQL syntax. To alleviate the gap between natural language and SQL queries, the Text-to-SQL task is proposed (Cai et al., 2018). Specifically, given a natural language question, Text-to-SQL aims to convert the question into a SQL query, making it more convenient for users to interact with database.
Traditional sequence-to-sequence methods mainly focus on fine-tune encoder-decoder models (Rai et al., 2023; Li et al., 2023a; Zeng et al., 2022), which are optimized for the encoding process and decoding process, including schema linking (Liu et al., 2021; Hui et al., 2021), building intermediate results between the natural language and SQL (Yu et al., 2018a; Wang et al., 2020; Gan et al., 2021), decoding constraints (Toutanova et al., 2021; Scholak et al., 2021; Suhr et al., 2020), etc. However, as the complexity of database information and user questions, these models encounter parameter constraints, which hamper their effectiveness and scalability.
@@
Recently, with the rise of LLMs, prompt-based Text-to-SQL approaches become more accessible. By simply providing prompts, LLMs can effortlessly get SQL queries and obtain outstanding performance (Liu et al., 2023; Chang and Fosler-Lussier, 2023; Dong et al., 2023; Pourreza and Rafiei, 2023; Gao et al., 2024). However, previous works in employing LLMs for Text-to-SQL prompting have three major limitations. (1) Inherent model bias: when generating SQL queries, models often exhibit inherent biases, such as GPT-3.5 is case-insensitive and may mistakenly convert an uppercase value in the question to a lowercase value in the SQL query (Liu et al., 2023). (2) Unexecutable SQL: although SQL is generated by the LLM, the LLM itself cannot determine whether the final SQL is executable or not. REACT (Yao et al., 2023) demonstrates that the LLM may generate illusory answers in question answering (QA) scenarios without external retrieval capabilities, incorporating retrieval into LLM generation enhances the accuracy and reliability of answer generation. Similarly, the LLM in the Text-to-SQL domain can also generate non-executable SQL queries. Therefore, the absence of external actions such as SQL execution limits the assurance of the executability of the generated SQL. (3) Expensive inference cost: recent works (Pourreza and Rafiei, 2023; Gao et al., 2024; Wang et al., 2023a) focus on leveraging GPT-4 for SQL generation. However, utilizing GPT-4 comes at a significant expense. Furthermore, most GPT-4 based methods (Pourreza and Rafiei, 2023; Gao et al., 2024; Wang et al., 2023a) rely on few-shot prompts, which leads to longer inputs and higher computational costs. For instance, DAIL-SQL (Gao et al., 2024) requires nearly \$0.1 per SQL query, MAC-SQL (Wang et al., 2023a) requires \$0.2 per SQL query, and DIN-SQL (Pourreza and Rafiei, 2023) requires nearly \$0.8 to generate each SQL query. If these methods are employed to generate thousands or even tens of thousands of SQL statements, it would be a catastrophic cost for ordinary users.
@@
To mitigate these issues, we introduce Semantic-Enhanced Text-to-SQL with Adaptive Refinement (SEA-SQL), an innovative framework based on zero-shot prompts and GPT-3.5 to accomplish the Text-to-SQL task. Our framework incorporates a semantic-enhanced schema to enrich database information and refine SQL queries through adaptive bias elimination and dynamic execution adjustment. Specifically, we augment the schema with question-related column values to enhance its semantic understanding. We then fine-tune and employ a bias eliminator to mitigate the inherent biases of the LLM during SQL generation. With dynamic execution adjustment, it ensures the executability of SQL queries by interleaving reflection and correction within the execution process.
@@
  • SEA-SQL significantly outperforms previous SOTA GPT-3.5 based Text-to-SQL method, and reaching an excellent performance competitive to GPT-4 with lower cost.
@@ 2 Related Work
With the rise of LLMs, individuals are progressively exploring the applications of LLMs, achieving successful outcomes particularly in the Text-to-SQL domain. For example, (Liu et al., 2023) conducts a comprehensive evaluation about zero-shot performance of GPT-3.5’s ability in Text-to-SQL task. C3 (Dong et al., 2023) proposes a zero-shot method utilizing GPT-3.5 to generate SQL, improving the generated SQL by providing clear layouts, calibrating bias prompts, and enhancing it through self-consistency (Wang et al., 2023b). DAIL-SQL (Gao et al., 2024) investigates various prompt representations and few-shot example selection methods to determine the most effective approach for efficient SQL generation. DIN-SQL (Pourreza and Rafiei, 2023) divides SQL generation into three stages: schema linking, SQL difficulty decomposition, and post-processing to refine SQL. MAC-SQL (Wang et al., 2023a) transforms the task of Text-to-SQL into a multi-agent collaboration task, constructs a selector for schema linking, a decomposer to break down questions and generate SQL, and a refiner to ensure executable SQL.
However, there are some issues with the methods mentioned earlier. Firstly, their ability to eliminate model bias is restricted to the model itself. This is achieved either by constraining model generation via prompts (Dong et al., 2023), or by creating rule-based prompts to correct SQL (Pourreza and Rafiei, 2023; Wang et al., 2023a). Since the model may not always adhere to prompts or recognize its own deficiencies, this strategy has limitations in eliminating model bias. Secondly, while MAC-SQL proposes ensuring SQL executability through execution, it identifies only five types of errors that the model may encounter in the prompt based on specific error cases. Consequently, the model’s effectiveness in addressing unknown errors is restricted, thereby limiting the model’s overall capability. Thirdly, most methods utilize GPT-4 for SQL generation, which comes with expensive costs. Although C3 (Dong et al., 2023) exclusively utilizes GPT-3.5, it suffers from the drawback of inferior performance.
Additionly, DAIL-SQL (Gao et al., 2024) presents a method to fine-tune open-source LLMs for the Text-to-SQL task. However, their performance is notably inferior, displaying a significant gap compared to other models in the GPT series.
@@ 3 Task Definition
Given a question $Q_{qes}$ and a database schema $S$ consisting of: (1) A set of tables, (2) A set of columns, and (3) A set of foreign keys. The Text-to-SQL task is to generate an SQL $A$ based on the question $Q_{qes}$ and the schema $S$. The generated SQL query $A$ should be executed on the database to provide the information expected by the question $Q_{qes}$.
@@ 4.1 Semantic-enhanced Schema
The study (Lin et al., 2020) points out the limitations of solely relying on table and column names, along with their foreign key relations, to grasp the semantics of a database schema with respect to a specific question. It is demonstrated that such an approach potentially overlook crucial contextual information contained within the actual data values of the database. Therefore, it becomes necessary to integrate the actual values present in the database into the SQL generation process to mitigate this limitation. While previous research (Lin et al., 2020) suggests augmenting the schema with values extracted from the columns mentioned in the question, this approach encounters challenges, particularly when the question does not explicitly specify the necessary values. For instance, consider a question like "the average salary of America", where the "country" column only contains the value "USA". In such cases, without supplemental information, it is difficult to accurately deduce the intended value format of the question.
To enable the LLM to understand the database comprehensively and utilize table names, column names, and value information correctly, we propose a simple yet effective method. Initially, we construct a BM25 index based on the database contents, leveraging the inherent structure and relationships within the data. Subsequently, we utilize this index to retrieve a subset of the top-k (e.g.,k=200) sample values from each column in the database. Building upon the "bridging" method introduced by (Lin et al., 2020), we select the top-m values from these top-k sets. We apply these sampled values to enhance the semantic representation of the database schema. As shown in Figure 2, compared to a standard schema, semantic-enhanced schema retrieves the top-m values of columns relevant to the question, assisting the LLM in better understanding the database content. This ensures that even in cases of ambiguous column names or particularly unique format column values, the LLM can still comprehend the database information accurately, enabling correct usage of table names, column names, and values.
The LLM then generates an SQL query based on both the original question $Q_{qes}$ and the database schema enhanced with semantic information $S$, following the zero-shot prompt format proposed by (Liu et al., 2023), as detailed in Appendix C:
@@ 4.2 Adaptive Bias Elimination
Due to the inherent bias of the LLM, it may generate SQL queries with specific model-related errors, e.g., GPT-3.5 prefer to use "LEFT JOIN" rather than "JOIN" (Liu et al., 2023; Dong et al., 2023). While approaches like (Pourreza and Rafiei, 2023) suggest using specific prompts to address biases, they are based on few-shot examples and manually add rules in prompts from the error case analysis, which may not cover all potential error cases and may not correct the errors effectively. To comprehensively eliminate the errors caused by the LLM inherent biases, we propose an adaptive bias eliminator, which adaptively discovers and corrects the inherent biases caused by the SQL generation.
$\bullet$ Objective. Previous study (Gao et al., 2024) indicates that a fine-tuned small LLM (e.g., Mistral-7B) does not yield comparable performance to GPT-3.5 or GPT-4 for directly generating SQL queries. The challenges encountered by the small LLM is attributed to the complexities involved in understanding SQL syntax and database information. In this paper, we propose a simplified task for the small LLM, suggesting it act as a bias eliminator of potential biases that might be introduced by the large SQL generation model (e.g., GPT-3.5). Therefore the small LLM does not need to fully comprehend all aspects of SQL syntax or complex database information. The primary role transition to identifying and eliminating inherent biases in the SQL generation model, ensures the production of high-quality and unbiased SQL queries.
Specifically, the bias eliminator receives two types of information: (1) SQL queries generated by the SQL generation model (pre-SQL), and (2) the execution results of SQL queries or the error messages generated by these queries. The former provides the coarse-grained information for the bias eliminator, while the latter offers a fine-grained information of the generated SQL queries, helping in the detection of potential biases in the generated SQL queries. The bias eliminator then outputs a revised SQL query, which removes any identified biases:
@@
where $A_{elim}$ denotes the bias-eliminated SQL, $M_{elim}$ represents the adaptive bias eliminator, and $r$ stands for the corresponding execution result of $A_{pre}$.
$\bullet$ Fine-tuning. To support the adaptive bias eliminator in identifying and mitigating biases introduced by the SQL generation LLM effectively, it must be exposed to a substantial number of bias cases produced by the SQL generation LLM. To obtain such training data, we conduct an extension to the standard training pairs: $(Q^{t}_{qes},A^{t}_{gold})$, where $Q^{t}_{qes}$ represents the question in the train set and $A^{t}_{gold}$ denotes the correct SQL query corresponding to the question $Q^{t}_{qes}$. Concretely, we use the SQL generation LLM to generate the pre-SQL $A^{t}_{pre}$ for each question, and obtain the execution result $r^{t}$ for each $A^{t}_{pre}$. The supplementary data $r^{t}$ provides the bias eliminator with a comprehensive understanding of where the SQL generation model is prone to making errors because of the inherent model biases. The new training pairs are expressed as: $(Q^{t}_{qes},S^{t},A^{t}_{pre},r^{t},A^{t}_{gold})$, where $S^{t}$ represents the semantic-enhanced schema in the train set.
Simultaneously, we expect that the adaptive bias eliminator possesses the ability to adjust SQL queries with model bias, while leaving SQL queries without model bias in their original format. To ensure that the bias eliminator does not modify bias-free pre-SQL queries which have different SQL formats compared to gold SQL queries but yield identical execution results, we adjust $A^{t}_{gold}$ as $A^{t}_{elim}$. If the pre-SQL query produces the same execution result as the gold SQL query, $A^{t}_{elim}$ is replaced with $A^{t}_{pre}$, otherwise, it retains $A^{t}_{gold}$.
Training objective for the bias eliminator $M_{elim}$ is to maximize the likelihood of output $A^{t}_{elim}$ across all $T$ training pairs:

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have significantly contributed to the progress of the Text-to-SQL task. A common requirement in many of these works is the post-correction of SQL queries. However, the majority of this process entails analyzing error cases to develop prompts with rules that eliminate model bias. And there is a weakness of execution verification for SQL queries. In addition, the prevalent techniques primarily depend on GPT-4 and few-shot prompts, resulting in expensive costs. To investigate the effective methods for SQL refinement in a cost-efficient manner, we introduce Semantic-Enhanced Text-to-SQL with Adaptive Refinement (SEA-SQL), which includes Adaptive Bias Elimination and Dynamic Execution Adjustment, aims to improve performance while minimizing resource expenditure with zero-shot prompts. Specifically, SEA-SQL employs a semantic-enhanced schema to augment database information and optimize SQL queries. During the SQL query generation, a fine-tuned adaptive bias eliminator is applied to mitigate inherent biases caused by the LLM. The dynamic execution adjustment is utilized to guarantee the executability of the bias eliminated SQL query. We conduct experiments on the Spider and BIRD datasets to demonstrate the effectiveness of this framework. The results demonstrate that SEA-SQL achieves state-of-the-art performance in the GPT3.5 scenario with 9%-58% of the generation cost. Furthermore, SEA-SQL is comparable to GPT-4 with only 0.9%-5.3% of the generation cost. Our code is available at https://github.com/545999961/SEA-SQL.
@@ 1 Introduction
Relational databases play a pivotal role in information technology by offering a structured and organized framework for data storage. Through tables, rows, and columns, data is efficiently organized, ensuring clarity and ease of management. However, users typically interact with databases through SQL queries, which is a barrier for those unfamiliar with SQL syntax. To alleviate the gap between natural language and SQL queries, the Text-to-SQL task is proposed Cai et al. (2018). Specifically, given a natural language question, Text-to-SQL aims to convert the question into a SQL query, making it more convenient for users to interact with database.
Traditional sequence-to-sequence methods mainly focus on fine-tuned encoder-decoder models Rai et al. (2023); Li et al. (2023b); Zeng et al. (2023), which are optimized for the encoding process and decoding process, including schema linking Liu et al. (2021); Hui et al. (2021), building intermediate results between the natural language and SQL Yu et al. (2018a); Wang et al. (2020); Gan et al. (2021), decoding constraints Yin and Neubig (2017); Scholak et al. (2021); Suhr et al. (2020), etc. As databases continue to expand in size, they often encompass a large number of tables and columns, complex inter-table relationships, diverse data types, and vast amounts of data. Consequently, user queries tend to exhibit complex structures, such as nested queries, aggregations, joins, or multiple conditions Li et al. (2023c). Both factors increase the complexity of inputs for the Text-to-SQL task. Sequence-to-sequence models have inherent limitations due to the small size of their parameters. As the complexity of input increases, these models may struggle to encode all necessary information efficiently, leading to performance bottlenecks.
@@
Recently, with the rise of LLMs, there has been a significant advancement in natural language understanding, contextual awareness, and scalability. LLMs are capable of effectively managing and interpreting complex queries and schema structures. This has made prompt-based Text-to-SQL approaches more accessible and effective. By simply providing prompts, LLMs can effortlessly get SQL queries and obtain outstanding performance Liu et al. (2023); Chang and Fosler-Lussier (2023); Dong et al. (2023); Pourreza and Rafiei (2023); Gao et al. (2024); Zhou et al. (2024). However, previous works in employing LLMs for Text-to-SQL prompting have three major limitations. (1) Inherent model bias: When generating SQL queries, models may exhibit biases due to patterns present in their training data. Such biases might manifest in specific preferences within the generated queries. For example, GPT-3.5 is case-insensitive and may inadvertently convert an uppercase value from the input into a lowercase value in the SQL query Liu et al. (2023). This behavior reflects the model’s inherent biases and can lead to inaccuracies in query formation. (2) Unexecutable SQL: Although SQL is generated by the LLM, they lack the capability to verify the executability of the final SQL query. When executed, these queries may encounter various errors, such as column mismatches or runtime errors. This limitation underscores the need for additional validation steps to ensure the generated SQL is functional and accurate. REACT Yao et al. (2023) demonstrates that the LLM may generate illusory answers in question answering (QA) scenarios without external retrieval capabilities, incorporating retrieval into LLM generation enhances the accuracy and reliability of answer generation. Similarly, the LLM in the Text-to-SQL domain can also generate non-executable SQL queries. Therefore, the absence of external actions such as SQL execution limits the assurance of the executability of the generated SQL. (3) Expensive inference cost: recent works Pourreza and Rafiei (2023); Gao et al. (2024); Wang et al. (2023a) focus on leveraging GPT-4 for SQL generation. However, utilizing GPT-4 comes at a significant expense. Furthermore, most GPT-4 based methods Pourreza and Rafiei (2023); Gao et al. (2024); Wang et al. (2023a) rely on few-shot prompts, which leads to longer inputs and higher computational costs. For instance, DAIL-SQL Gao et al. (2024) requires nearly \$0.1 per SQL query, MAC-SQL Wang et al. (2023a) requires \$0.2 per SQL query, and DIN-SQL Pourreza and Rafiei (2023) requires nearly \$0.8 to generate each SQL query. If these methods are employed to generate thousands or even tens of thousands of SQL statements, it would be a catastrophic cost for ordinary users.
@@
To mitigate these issues, we introduce Semantic-Enhanced Text-to-SQL with Adaptive Refinement (SEA-SQL), an innovative framework based on zero-shot prompts and GPT-3.5 to accomplish the Text-to-SQL task in an economical manner. Our framework incorporates a semantic-enhanced schema to enrich database information and refine SQL queries through adaptive bias elimination and dynamic execution adjustment. Specifically, we augment the schema with question-related column values to enhance its semantic understanding. We then fine-tune and employ a bias eliminator to mitigate the inherent biases of the LLM during SQL generation. With dynamic execution adjustment, it ensures the executability of SQL queries by interleaving reflection and correction within the execution process.
@@
  • SEA-SQL significantly outperforms previous SOTA GPT-3.5 based Text-to-SQL method, and obtains an excellent performance competitive to GPT-4 with lower cost.
@@ 2 Related Work
With the rise of LLMs, individuals are progressively exploring the applications of LLMs, achieving successful outcomes particularly in the Text-to-SQL domain. For example, ChatGPT’s Evaluation Liu et al. (2023) conducts a comprehensive evaluation about zero-shot performance of GPT-3.5’s ability in Text-to-SQL task. C3 Dong et al. (2023) proposes a zero-shot method utilizing GPT-3.5 to generate SQL, improving the generated SQL by providing clear layouts, calibrating bias prompts, and enhancing it through self-consistency Wang et al. (2023c). DAIL-SQL Gao et al. (2024) investigates various prompt representations and few-shot example selection methods to determine the most effective approach for efficient SQL generation. DIN-SQL Pourreza and Rafiei (2023) divides SQL generation into three stages: schema linking, SQL difficulty decomposition, and post-processing to refine SQL. MAC-SQL Wang et al. (2023a) transforms the task of Text-to-SQL into a multi-agent collaboration task, constructs a selector for schema linking, a decomposer to break down questions and generate SQL, and a refiner to ensure executable SQL.
However, there are some issues with the methods mentioned earlier. Firstly, their ability to eliminate model bias is restricted to the model itself. This is achieved either by constraining model generation via prompts Dong et al. (2023), or by creating rule-based prompts to correct SQL Pourreza and Rafiei (2023); Wang et al. (2023a). Since the model may not always adhere to prompts or recognize its own deficiencies, this strategy has limitations in eliminating model bias. Secondly, while MAC-SQL proposes ensuring SQL executability through execution, it identifies only five types of errors that the model may encounter in the prompt based on specific error cases. Consequently, the model’s effectiveness in addressing unknown errors is restricted, thereby limiting the model’s overall capability. Thirdly, most methods utilize GPT-4 for SQL generation, which comes with expensive costs. Although C3 Dong et al. (2023) exclusively utilizes GPT-3.5, it suffers from the drawback of inferior performance.
Additionly, DAIL-SQL Gao et al. (2024) presents a method to fine-tune open-source LLMs for the Text-to-SQL task. However, their performance is notably inferior, displaying a significant gap compared to other models in the GPT series.
@@ 3 Task Definition
Given a question $Q_{qes}$ and a database schema $S$ consisting of:
  1. 1.
    A set of tables $T={t_{1},t_{2},\dots,t_{N}}$, where $N$ is the total number of tables. Each $t_{i}$ represents the $i$-th table.
  2. 2.
    A set of columns $C={c^{1}_{1},\dots,c^{1}_{n_{1}},\dots,c^{N}_{1},\dots,}$ ${c^{N}_{n_{N}}}$, where $n_{i}$ denotes the number of columns in table $t_{i}$, and $c^{i}_{j}$ represents the $j$-th column of table $t_{i}$.
  3. 3.
    A set of values $V={v^{1}_{1},\dots,v^{1}_{m},\dots,v^{M}_{1},\dots,v^{M}_{m}}$, where $M=\sum_{i=1}^{N}n_{i}$ is the total number of columns. Each column has $m$ sample values, and $v^{j}_{k}$ denotes the $k$-th value of column $c^{j}$.
  4. 4.
    A set of foreign keys $R=\left\{\left(t_{i}.c_{k}=t_{j}.c_{l}\right)\right\}$, where $t_{i},t_{j}\in T,c_{k},c_{l}\in C$.
The Text-to-SQL task is to generate an SQL $A$ based on the question $Q_{qes}$ and the schema $S$. This generated SQL can be executed on the database to get the answer of the question $Q_{qes}$.
@@ 4.1 Semantic-enhanced Schema
The study Lin et al. (2020) points out the limitations of solely relying on table and column names, along with their foreign key relations, to grasp the semantics of a database schema with respect to a specific question. It is demonstrated that such an approach potentially overlooks crucial contextual information contained within the actual data values of the database. Therefore, it becomes necessary to integrate the actual values present in the database into the SQL generation process to mitigate this limitation. While previous research Lin et al. (2020) suggests augmenting the schema with values extracted from the columns mentioned in the question, this approach encounters challenges, particularly when the question does not explicitly specify the necessary values. For instance, consider a question like "the average salary of America", where the "country" column only contains the value "USA". In such cases, without supplemental information, it is difficult to accurately deduce the intended value format of the question.
To enable the LLM to understand the database comprehensively and utilize table names, column names, and value information correctly, we propose a simple yet effective method Wan and Han (2024). Although the "bridging" method Lin et al. (2020) yields excellent performance, its application is limited by slow processing speeds when handling large datasets. Additionally, it extracts only the values explicitly mentioned in the question, which can result in insufficient information capture. To address these limitations, we propose a new retrieval paradigm that combines BM25 Li et al. (2024) with the "bridging" method Wang et al. (2023b). In this integrated approach, "bridging" functions as a second-stage retrieval mechanism, enhancing the overall efficiency and depth of information retrieval.
Initially, we construct a BM25 index for each column of the database, leveraging the inherent structure and relationships within the data. Using this index, we retrieve a subset of the top-k sample values (e.g., k=1000) from each column in the database. We then apply the "bridging" method Lin et al. (2020) for string matching. If any of the top-k values perfectly match specific words in a given question, these values are moved to the new top-1 position. Finally, we select the top-m values (e.g., m=2) from these rearranged lists as sampled values. We apply these sampled values to enhance the semantic representation of the database schema. As shown in Figure 2, compared to a standard schema, semantic-enhanced schema retrieves the top-m values of columns relevant to the question, assisting the LLM in better understanding the database content. This ensures that even in cases of ambiguous column names or particularly unique format column values, the LLM can still comprehend the database information accurately, enabling correct usage of table names, column names, and values.
The LLM then generates an SQL query based on both the original question $Q_{qes}$ and the database schema enhanced with semantic information $S$, following the zero-shot prompt format proposed by Liu et al. (2023), as detailed in Appendix C:
@@ 4.2 Adaptive Bias Elimination
Due to inherent biases, LLMs may generate SQL queries that contain specific model-related errors. These errors often occur because the models tend to favor certain patterns influenced by biases present in the training data, e.g., GPT-3.5 prefer to use "LEFT JOIN" rather than "JOIN" Liu et al. (2023); Dong et al. (2023). While approaches like DIN-SQL Pourreza and Rafiei (2023) suggest using specific prompts to address biases, they are based on few-shot examples and manually add rules in prompts from the error case analysis, which may not cover all potential error cases and may not correct the errors effectively. To comprehensively eliminate the errors caused by the LLM inherent biases, we propose an adaptive bias eliminator, which adaptively discovers and corrects the inherent biases caused by the SQL generation.
$\bullet$ Objective. Previous study Gao et al. (2024) indicates that a fine-tuned small LLM (e.g., Mistral-7B) does not yield comparable performance to GPT-3.5 or GPT-4 for directly generating SQL queries. The challenges encountered by the small LLM is attributed to the complexities involved in understanding SQL syntax and database information. In this paper, we propose a simplified task for the small LLM, suggesting it act as a bias eliminator of potential biases that might be introduced by the large SQL generation model (e.g., GPT-3.5). Therefore the small LLM does not need to fully comprehend all aspects of SQL syntax or complex database information. The primary role transition to identifying and eliminating inherent biases in the SQL generation model, ensures the production of high-quality and unbiased SQL queries.
Specifically, the bias eliminator receives two types of information: (1) SQL queries generated by the SQL generation model (pre-SQL), and (2) the execution results of queries include the execution information (success or failure) and their corresponding results. The former provides the coarse-grained information for the bias eliminator, while the latter offers a fine-grained information of the generated SQL queries, helping in the detection of potential biases in the generated SQL queries. The bias eliminator then outputs a revised SQL query, which removes any identified biases:
@@
where $A_{elim}$ denotes the bias-eliminated SQL, $M_{elim}$ represents the adaptive bias eliminator, and $r$ stands for the corresponding execution result of $A_{pre}$. If an error occurs during execution and no results are obtained, the content of $r$ will be the error message.
Figure 3: The fine-tuning process of Adaptive Bias Elimination. A checkmark () indicates that the execution result of $A^{t}_{\text{pre}}$ is identical to the execution result of $A^{t}_{\text{gold}}$, while a crossmark () indicates they differ.
$\bullet$ Fine-tuning. As shown in Figure 3, to support the adaptive bias eliminator in identifying and mitigating biases introduced by the SQL generation LLM effectively, it must be exposed to a substantial number of bias cases produced by the SQL generation LLM. To obtain such training data, we conduct an extension to the standard training pairs: $(Q^{t}_{qes},S^{t},A^{t}_{gold})$, where $t$ indicates the training data, $Q^{t}_{qes}$ represents the question in the train set and $A^{t}_{gold}$ denotes the correct SQL query corresponding to the question $Q^{t}_{qes}$ and $S^{t}$ represents the semantic-enhanced schema in the train set. Concretely, we use the SQL generation LLM to generate the pre-SQL $A^{t}_{pre}$ for each question, and obtain the execution result $r^{t}$ for each $A^{t}_{pre}$. The supplementary data $r^{t}$ provides the bias eliminator with a comprehensive understanding of where the SQL generation model is prone to making errors because of the inherent model biases.
We expect the adaptive bias eliminator to be capable of modifying SQL queries that contain model bias, while preserving the original format of queries that are free from such bias. If a pre-SQL query yields execution results identical to those of a gold SQL query, then the bias eliminator should leave the pre-SQL query unchanged, denoted as $A^{t}_{elim}=A^{t}_{pre}$. Conversely, if the pre-SQL query execution results deviate from those of the gold SQL query, the bias eliminator should adjust the query to match the gold, reflected by $A^{t}_{elim}=A^{t}_{gold}$. This ensures that modifications are only made when necessary and that the integrity of bias-free queries is maintained. Finally, the new training pairs are expressed as: $(Q^{t}_{qes},S^{t},A^{t}_{pre},r^{t},A^{t}_{elim})$.
The training objective for the small LLM bias eliminator, denoted as $M_{elim}$ is to maximize the likelihood of output $A^{t}_{elim}$ across all $T$ training pairs: